Francesca would like to lease a coffee cart in Whistler, B.C. The lease is $800 per month, and a city licence to sell food and beverages costs $20 per month. The lessor of the stand has shown Francesca records indicating that gross revenues average $32 per hour. The out-of-pocket costs for ingredients are generally about 40% of gross revenues. Last year she paid 25% of her income in taxes.

Francesca pays $1,000 per month for her condominium. She could store the cart overnight in the condo’s garage, which is currently unused. Real estate developers in Whistler estimate that about 20% of the cost of a residential building is for the garage.

At present, Francesca is earning $2,400 per month as a ski instructor for one of the big ski areas. In the summertime, she earns about the same income by working as a kayaking instructor.

1. For each piece of quantitative information in this problem, indicate whether it is relevant to Francesca’s decision.

Quantitative information
1. Cart lease $800 per month:

Relevant or Irrelevant

2. City license $20 per month:

Relevant or Irrelevant

3. Ingredients 40% of revenue: Relevant or Irrelevant
4. Last year’s income tax rate of 25%:

Relevant or Irrelevant

5. Condo rent of $1,000 per month and 20% of condo cost for garage:

Relevant or Irrelevant

6. Current income $2,400 per month:

Relevant or Irrelevant

2.

If Pearl leases the cart and works 30 days in a month, how many hours will she have to work
each day, on average, to be at least as well off financially as she is in her current job?
each day, on average, to be at least as well off financially as she is in her current job?

3.

If Pearl wants to work only 25 days per month, how much will revenues have to increase PER HOUR for her to work 4 hours per day and be as finically well off as she is in her current job?
Answer #1

12 1. For each piece of quantitative information in this problem, indicate whether it is relevant to Francescas decision. QuThe monthly cost function is estimated as: TC = 820 +0.40* Revenue 2,400 = (32-0.40*32) *Hours per month - 820 2,400 +820 = (3 If Pearl wants to work only 25 days per month, how much will revenues have to increase PER HOUR for her to won Total hours